OpenStudy (anonymous):

Compare and Contrast: Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution. Equation #1 Equation #2 6 - x = 5x + 30 4x - 8 = 3x + 2 The solution to equation #1 is larger than the solution to equation #2. None of the statements above describes solutions to equations shown. The solution to equation #1 is smaller than the solution to equation #2. The solution to equation #1 is the same as the solution to equation #2.

OpenStudy (anonymous):

The solution to equation #1 is smaller than the solution to equation #2.

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

eq #1: 6 - x = 5x + 30 6 = 4x+ 30 4x = -24 x = -6 eq #2: 4x - 8 = 3x + 2 x = 6